# Uncertainty-Guided Coarse-to-Fine Tumor Segmentation with Anatomy-Aware Post-Processing

### v1

@@ Abstract
uncertainty, tumor segmentation, lung cancer,
@@ I Introduction
Our Contribution and Findings. To address these needs, we propose a two-stage, end-to-end segmentation framework that combines coarse tumor localization with refined region-of-interest (ROI) segmentation. In the first stage, a full-volume model identifies candidate tumor regions. These are refined through a clinically informed post-processing pipeline that applies anatomical constraints, including lung overlap thresholds, minimum surface distance, and region-specific filters. Rather than introducing an explicit localization model, we treat segmentation as a surrogate for localization. We show that lightweight, anatomy-aware post-processing can yield high spatial fidelity without increasing computational load.
@@
Figure 1: Overview of the proposed two-stage tumor segmentation framework. The first-stage model performs coarse segmentation on the full CT scan and outputs both a tumor prediction and an uncertainty map. Post-processing—including edge enhancement, lung masking, and ROI cropping—uses these outputs to generate a refined input region. The second-stage model conducts fine segmentation on the cropped ROI to produce the final tumor mask. The uncertainty map enhances interpretability and supports adaptive refinement around ambiguous boundaries.
Figure 2: Example patient cases from our proposed two-stage segmentation pipeline. From left to right: full CT image, ground truth mask, initial prediction, post-processed prediction, and the region-of-interest (ROI) prediction with uncertainty map overlay. Our pipeline demonstrates improved tumor delineation and spatial accuracy through uncertainty modeling and anatomical post-processing.
@@ III Methodology
This study proposes a two-stage, end-to-end segmentation framework that integrates coarse tumor localization, uncertainty-aware ROI refinement, and anatomically guided post-processing. The goal is to enhance segmentation precision, robustness, and interpretability in thoracic CT scans while maintaining computational efficiency. The methodology is structured to evaluate segmentation performance under variations in ROI context, component selection, uncertainty modeling, model architecture, and dataset generalization.
@@ III-D Uncertainty-Aware Loss Functions
where $U(x)$ represents the uncertainty at pixel $x$, computed from the variance of multiple stochastic forward passes using Monte Carlo (MC) Dropout. To support this, dropout layers remain active during inference, enabling the generation of variance-based uncertainty maps that inform both loss reweighting during training and interpretability during evaluation.
@@ III-E Model Architectures
To evaluate the sensitivity of the proposed framework to architectural variations, multiple 3D segmentation backbones–UNet, UNETR, and Swin UNETR—were integrated into the end-to-end pipeline. A baseline model was trained on full-resolution CT volumes using a single-stage approach, serving as a reference for comparison.
In contrast, the ROI-based segmentation models were trained on region-of-interest (ROI) crops extracted using either tight bounding (0-margin) or padded context (16-margin). Each model was evaluated under different loss configurations, including both standard and uncertainty-aware objectives.
@@ III-F Post-Processing
Post-processing was applied exclusively to the outputs of the first-stage coarse segmentation model to refine tumor candidate regions before ROI extraction. This stage ensures that only anatomically plausible and spatially relevant regions are passed to the second-stage model, reducing false positives, improving ROI focus, and contributing to lower predictive uncertainty by filtering out ambiguous or anatomically inconsistent candidates. The refinement process involved a series of spatial filters informed by clinical and anatomical knowledge, which we review in the following:
@@
    Top-K Component Selection (Optional). After filtering, the $K$ largest components by voxel volume were optionally retained to reduce clutter and improve focus. This can be interpreted as a form of spatial attention, guiding the second-stage model to prioritize the most prominent tumor candidates while suppressing ambiguous or minor regions. These post-processing heuristics reflect common clinical scenarios. While most gross tumor volumes (GTVs) lie within the lung parenchyma and are well-captured by lung masks, tumors may also extend into extrapulmonary areas such as the pleura or mediastinum. The pipeline is therefore designed not only to suppress anatomically implausible predictions but also to preserve clinically relevant yet spatially borderline tumors that may extend beyond lung boundaries. By filtering for structural plausibility and emphasizing dominant components, the post-processing step enhances the quality of ROI extraction and reduces noise in the downstream fine segmentation stage.
@@ IV Results and Comparative Analysis
Table I: Comparison of the original model vs. the proposed end-to-end system using Top-1 ROI. We compare two ROI models trained on different settings (ROI-0margin and ROI-16margin).
@@
End-to-End (Top-1, ROI-16margin model) 0.5917 10.79
End-to-End (Top-1, ROI-0margin model) 0.6447 8.89
Table II: Effect of the number of components used during inference. We evaluate the end-to-end system performance using Top-1, Top-2, and all valid components after post-processing.
@@
Table IV: Effect of using different lung masks in the end-to-end system. We compare the original lung mask with a newly generated one.
@@
Table V: Impact of using an uncertainty-aware loss function in the ROI model. All models are evaluated in an end-to-end setup using Top-1 ROI.
@@
Two-Stage vs. Baseline. The two-stage framework was first compared to a baseline single-stage full-volume segmentation model. As shown in Table I, both ROI-0margin and ROI-16margin configurations outperformed the baseline. The ROI-0margin setup achieved the highest Dice score (0.6447) and lowest HD95 (8.89 mm), indicating that tighter cropping yields more accurate segmentation.
Impact of Component Selection. Table II reports the results on the impact of component selection. We found that retaining only the largest component (Top-1) led to the best results, and including additional components (Top-2 or All Valid) reduced Dice and increased HD95, suggesting that false positives and irrelevant regions impair segmentation.
@@
Summary of Key Findings. The proposed two-stage segmentation framework significantly outperformed the baseline single-stage approach, particularly when using tightly cropped ROI inputs. The use of Top-1 component selection led to higher accuracy and reduced boundary errors, while including additional components introduced false positives and degraded performance. A clear negative correlation was observed between the number of components and segmentation quality, highlighting the value of post-processing. Moreover, lung mask quality had little effect, affirming the reliability of automated segmentation. While standard Dice Cross-Entropy loss yielded the best metrics, uncertainty-aware objectives enhanced spatial calibration and interpretability, key for clinical use. Swin UNETR showed the greatest improvement on the Orlando dataset, with more modest gains on NSCLC, indicating that initial model quality affects but does not fully determine refinement success. Overall, the results highlight the complementary value of ROI refinement, component filtering, and uncertainty modeling for accurate, clinically meaningful tumor segmentation.
@@ V Conclusion
This study demonstrates the effectiveness of a two-stage, anatomy-aware segmentation framework for tumor delineation in lung CT. By combining ROI refinement, component filtering, and post-processing heuristics, the proposed approach significantly improves segmentation accuracy and reliability. Cases with fewer connected components exhibited higher Dice scores and lower HD95 values, with statistically significant correlations confirming the relationship between component reduction and performance improvement.
Across both the Orlando Health and NSCLC datasets, ROI cropping and Top-1 component selection consistently outperformed baseline full-volume models. Although uncertainty-aware loss functions did not always yield higher Dice or HD95 scores, they enhanced spatial calibration and boundary consistency which are critical factors in clinical interpretation.
The integration of anatomical priors, lightweight postprocessing, and uncertainty modeling yields a robust and interpretable segmentation pipeline. These findings support the use of modular, post-hoc strategies to refine predictions from modern segmentation networks, especially in scenarios with limited supervision or heterogeneous imaging conditions.

### v2

@@ Abstract
uncertainty, tumor segmentation, lung cancer
@@ I Introduction
Our Contribution and Findings. To address these needs, we propose a two-stage, fully automated segmentation framework that combines coarse tumor localization with refined region-of-interest (ROI) segmentation. In the first stage, a full-volume model identifies candidate tumor regions. These are refined through a clinically informed post-processing pipeline that applies anatomical constraints, including lung overlap thresholds, minimum surface distance, and region-specific filters. Rather than introducing an explicit localization model, we treat segmentation as a surrogate for localization. We show that lightweight, anatomy-aware post-processing can yield high spatial fidelity without increasing computational load.
@@
Figure 1: Overview of the proposed two-stage tumor segmentation framework. The first-stage model performs coarse segmentation on the full CT scan and outputs both a tumor prediction and an uncertainty map. Post-processing module uses these outputs to generate a refined input region. The second-stage model conducts fine segmentation on the cropped ROI to produce the final tumor mask. The uncertainty map enhances interpretability and supports adaptive refinement around ambiguous boundaries.
Figure 2: Example patient cases from our two-stage segmentation pipeline. From left to right: full CT image, ground truth mask, initial prediction, post-processed prediction, and the region-of-interest prediction with uncertainty map overlay. Our pipeline demonstrates improved tumor delineation and spatial accuracy through uncertainty modeling and anatomical post-processing.
@@ III Methodology
This study proposes a two-stage, fully automated segmentation framework that integrates coarse tumor localization, uncertainty-aware ROI refinement, and anatomically guided post-processing. The goal is to enhance segmentation precision, robustness, and interpretability in thoracic CT scans while maintaining computational efficiency. The methodology is structured to evaluate segmentation performance under variations in ROI context, component selection, uncertainty modeling, model architecture, and dataset generalization.
@@ III-D Uncertainty-Aware Loss Functions
where $U(x)$ represents the uncertainty at pixel $x$, computed from the variance of multiple forward passes using Monte Carlo (MC) Dropout. To support this, dropout layers remain active during inference, enabling the generation of variance-based uncertainty maps that inform both loss reweighting during training and interpretability during evaluation.
@@ III-E Model Architectures
To assess the framework’s sensitivity to architectural variations, we integrated three 3D segmentation backbones–UNet, UNETR, and Swin UNETR–into the automated pipeline. A baseline single-stage model trained on full-resolution CT volumes served as a reference. ROI-based models were trained on crops extracted using either tight (0-margin) or padded (16-margin) bounding boxes. All models were evaluated under both standard and uncertainty-aware loss functions.
@@ III-F Post-Processing
Post-processing was applied solely to the first-stage coarse segmentation outputs to refine tumor candidate regions prior to ROI extraction. This step ensured that only anatomically plausible and spatially relevant regions were forwarded to the second-stage model, reducing false positives, enhancing ROI precision, and lowering predictive uncertainty by eliminating ambiguous or anatomically inconsistent candidates. The refinement process employed spatial filters guided by clinical and anatomical knowledge, as detailed below.
@@
    Top-K Component Selection (Optional). After filtering, the top $K$ largest components by voxel volume were optionally retained to reduce clutter and enhance focus—a form of spatial attention that guides the second-stage model toward prominent tumor candidates while suppressing minor or ambiguous regions. These heuristics reflect clinical realities: while most gross tumor volumes (GTVs) reside within the lung parenchyma, some extend into extrapulmonary regions like the pleura or mediastinum. The pipeline is thus designed to suppress anatomically implausible predictions while preserving borderline but clinically relevant tumors. This structural filtering improves ROI extraction and reduces noise in downstream fine segmentation.
@@ IV Results and Comparative Analysis
Table I: Comparison of the original model vs. the proposed fully automated system using Top-1 ROI. We compare two ROI models trained on different settings (ROI-0margin and ROI-16margin).
@@
fully automated (Top-1, ROI-16margin model) 0.5917 10.79
fully automated (Top-1, ROI-0margin model) 0.6447 8.89
Table II: Effect of the number of components used during inference. We evaluate the fully automated system performance using Top-1, Top-2, and all valid components after post-processing.
@@
Table IV: Effect of using different lung masks in the fully automated system. We compare the original lung mask with a newly generated one.
@@
Table V: Impact of using an uncertainty-aware loss function in the ROI model. All models are evaluated in an fully automated setup using Top-1 ROI.
@@
Two-Stage vs. Baseline. The two-stage framework was first compared to a baseline single-stage full-volume segmentation model. As shown in Table I, both ROI-0margin and ROI-16margin configurations outperformed the baseline. The ROI-0margin setup achieved the highest Dice score (0.6447) and lowest HD95 (8.89 mm), indicating that tighter cropping yields more accurate segmentation. Here, the “Original Full CT” model refers to the baseline model trained on full-resolution CT volumes using the Swin UNETR architecture, without any post-processing, ROI refinement, or uncertainty modeling. The “Dice (Init)” values reported in Table I denote the performance of each architecture before applying the proposed two-stage cascade and loss refinements. All baseline models were trained using the same optimizer, augmentations, and learning schedule as their refined counterparts to ensure fair comparison.
Impact of Component Selection. Table II reports the results on the impact of component selection. We found that retaining only the largest component (Top-1) led to the best results, and including additional components (Top-2 or All Valid) reduced Dice and increased HD95, suggesting that false positives and irrelevant regions impair segmentation. However, this gain in precision may come at the expense of reduced sensitivity to multifocal tumors or small satellite lesions. In cases where detecting all foci is essential, retaining multiple components may be preferable, even if it slightly lowers the average Dice score. To suppress noise and fragmented predictions, our post-processing pipeline discards components with 50 or fewer voxels, without imposing an upper size limit. Sensitivity analysis across thresholds from 50 to 150 voxels showed Dice variations of less than 0.002, but thresholds above 150 began excluding true tumor regions. These results support a conservative lower-bound threshold to balance anatomical plausibility and segmentation integrity.
@@
Summary. The proposed two-stage segmentation framework outperformed the baseline single-stage approach, especially with tightly cropped ROI inputs. Top-1 component selection improved accuracy and reduced boundary errors, while including additional components increased false positives and degraded performance. A strong negative correlation between component count and segmentation quality underscored the importance of post-processing. Lung mask quality had minimal impact, confirming the robustness of automated segmentation. While standard Dice Cross-Entropy loss achieved the highest metrics, uncertainty-aware objectives improved spatial calibration and interpretability-critical for clinical use. Swin UNETR yielded the greatest gains on the Orlando dataset, with smaller improvements on NSCLC, suggesting that initial model quality influences but does not limit refinement. Overall, the results emphasize the complementary benefits of ROI refinement, component filtering, and uncertainty modeling for accurate and clinically meaningful tumor segmentation.
@@ V Conclusion
This study demonstrates the effectiveness of a two-stage, anatomy-aware segmentation for lung tumor delineation in CT scans. By integrating ROI refinement, component filtering, and heuristic post-processing, the approach significantly enhances segmentation accuracy and reliability. Fewer connected components were associated with higher Dice scores and lower HD95 values, with statistically significant correlations supporting this relationship. Across both datasets, ROI cropping and Top-1 component selection consistently outperformed full-volume baselines. While uncertainty-aware losses did not always improve Dice or HD95 scores, they enhanced spatial calibration and boundary consistency.
Our pipeline consistently improved performance across both datasets, though the extent varied with dataset characteristics. The Orlando Health dataset saw greater benefit from post-processing due to small, isolated tumors well-suited to component filtering. In contrast, the NSCLC dataset, with larger or multifocal lesions, was more sensitive to ROI and component thresholds. Despite potential concerns about the generalizability of heuristic post-processing, the framework remained stable under modest parameter variations. Sensitivity analysis confirmed that changes in lung overlap and voxel size thresholds had minimal impact, indicating robustness with minimal tuning. Initial hyperparameters were guided by clinical priors, including typical tumor sizes and lung involvement patterns reported in NSCLC literature.
Integrating anatomical priors, lightweight post-processing, and uncertainty modeling results in a robust and interpretable segmentation pipeline. These findings support modular, post-hoc refinement strategies as effective complements to modern segmentation networks, particularly under limited supervision or heterogeneous imaging conditions.